OpenStudy (anonymous):

General form of a parabola vertex- (4,1) x-intercepts (3,0) and (5,0)

OpenStudy (anonymous):

ok... do you know the vertex form for the equation for a parabola???

OpenStudy (anonymous):

y=a(x-h)^2+k

OpenStudy (anonymous):

ok... good so we have: \(\large y=a(x-4)^2+1 \) since we know the parabola passes through the given points, plug those values in so we can find "a"....

OpenStudy (anonymous):

All I know is that the vertex is (h,k) so y=a(x-4)^2+1 I just don't know how to find a

OpenStudy (anonymous):

i'm a mind reader!!!

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

ok... let's plug in (3,0): \(\large 0=a(3-4)^2+1 \) solve for a.... wat u got???

OpenStudy (anonymous):

0=0?

OpenStudy (anonymous):

not quite... try again...

OpenStudy (anonymous):

\(\large 0=a(3-4)^2+1 \) \(\large 0=a(-1)^2+1 \) \(\large 0=a\cdot 1+1 \) \(\large 0=a+1 \) a = ???

OpenStudy (anonymous):

Omg, I'm retarded lol -1=a

OpenStudy (anonymous):

ok... just to make sure, let's see if we get the same value for a if we plugged in the other point (5, 0)... do we get the same value???

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

good, so the VERTEX form is: \(\large y=-(x-4)^2+1 \) you want GENERAL form so what we need to do is expand the right side and move the y over to the right side.....

OpenStudy (anonymous):

So the equation would be y=-1(x-4)^2+1

OpenStudy (anonymous):

mind reader!!!

OpenStudy (anonymous):

Lol, it's perfect because your owl looks shocked

OpenStudy (anonymous):

haha...:) ok... expand that right side....

OpenStudy (anonymous):

0=-y-(x-4)^2+1

OpenStudy (anonymous):

ok.. general form looks like this: \(\large 0=Ax^2+Bx+Cy+D \) so you need to expand (x-4)^2

OpenStudy (anonymous):

(x-4)^2=x^2-8x+16

OpenStudy (anonymous):

ok... so we still have: \(\large 0= -y-(x^2-8x+16)+1\) so arranging and simplifying them we have: \(\large 0= -y-x^2+8x-16+1\) \(\large 0= -x^2+8x-y-15\) OR \(\large x^2-8x+y+15=0\)

OpenStudy (anonymous):

So that's the answer? x^2-8x+y+15=0

OpenStudy (anonymous):

either answer will work... they're both in general form....

OpenStudy (anonymous):

yes... :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw....:)